OpenStudy (anonymous):

Rewrite the expression ln(a+b)+5ln(a−b)−4ln c as a single logarithm lnA. Then the function A=

OpenStudy (fibonaccichick666):

so what properties do you know about logs?

OpenStudy (anonymous):

would it be that ln(ab)=ln(a+b) and that ln(a/b)=ln(a-b)

OpenStudy (anonymous):

that's the tricky part. what you said is not correct. ln(a) + ln(b) = ln(ab)

OpenStudy (anonymous):

i have the answer ready to paste whenever you want it. let me know when you're done attempting ^_^

OpenStudy (fibonaccichick666):

also remember that you have another property that deals with exponents and I agree with @Euler271 's correction.

OpenStudy (anonymous):

oh ok im not sure if its correct but i got ln(ab)+ln(a/b)^5-ln(c)^4 is that how you do it

OpenStudy (fibonaccichick666):

take @Euler271 's coorection and use that property instead of the one you typed because it is incorrect .

OpenStudy (anonymous):

ill paste my answer. let me know if you want me to clarify any part of it: the answer i get is: \[\ln \frac{ (a+b)(a-b)^{5} }{ c^{4} }\]

OpenStudy (fibonaccichick666):

^this is correct^

OpenStudy (fibonaccichick666):

you did do the exponents correctly though :)

OpenStudy (anonymous):

i should have put brackets around whole thing: \[\ln (\frac{ (a+b)(a-b)^{5} }{ c^{4} })\]

OpenStudy (fibonaccichick666):

haha yes you should have. 10 pts from Ravenclaw lol

OpenStudy (anonymous):

oh ok i think i understand